# Jingxuan Li UCLA

### v2

@@ Abstract
The recent progress in Vision-Language Models (VLMs) has broadened the scope of multimodal applications. However, evaluations often remain limited to functional tasks, neglecting abstract dimensions such as personality traits and human values. To address this gap, we introduce Value-Spectrum, a novel Visual Question Answering (VQA) benchmark aimed at assessing VLMs based on Schwartz’s value dimensions that capture core values guiding people’s preferences and actions. We designed a VLM agent pipeline to simulate video browsing and constructed a vector database comprising over 50,000 short videos from TikTok, YouTube Shorts, and Instagram Reels. These videos span multiple months and cover diverse topics, including family, health, hobbies, society, technology, etc. Benchmarking on Value-Spectrum highlights notable variations in how VLMs handle value-oriented content. Beyond identifying VLMs’ intrinsic preferences, we also explored the ability of VLM agents to adopt specific personas when explicitly prompted, revealing insights into the adaptability of the model in role-playing scenarios. These findings highlight the potential of Value-Spectrum as a comprehensive evaluation set for tracking VLM alignments in value-based tasks and abilities to simulate diverse personas.
@@
Jingxuan Li1,*, Yuning Yang1,*, Shengqi Yang2, Linfan Zhang1, Ying Nian Wu1, 1University of California, Los Angeles, 2Los Alamos National Laboratory, Correspondence: madili@g.ucla.edu
@@ 1 Introduction
Figure 1: Exploring Value-Driven Role-Playing in Vision-Language Models. This study investigates how VLMs adopt assigned personas to align value traits and preferences within social media contexts.
@@
Figure 2: Overview of short video screenshots distribution of Value-Spectrum Dataset. We collected an abundance of short video screenshots relevant to 10 Schwartz values. The area of circles centered at each keyword represents the number of relevant videos in the database.
@@
Vision-Language Models (VLMs), built upon Large Language Models (LLMs) with pre-trained vision encoders through cross-modal alignment training, have shown impressive perceptual and cognitive capabilities in tasks like VQA and image captioning Zhou et al. (2019); Radford et al. (2021); Zhang et al. (2023); Lu et al. (2024). Recent research has identified that LLMs exhibit distinct preferences Li et al. (2024), personalities Serapio-García et al. (2023), and values Ren et al. (2024). In addition, some studies have explored the potential of LLMs as role-playing agents to simulate various personas Wang et al. (2023b); Chen et al. (2024a). Questions thus arise about whether VLMs, as visual extensions of LLMs, also exhibit inherent preferences and whether they can be induced to role-play specific personas.
@@
To address these concerns, our study explores two key questions: (1) Do VLMs exhibit preference traits? (2) Could VLMs adapt their traits to role-play specific human-designed personas, aligning their behaviors and preferences to match predefined roles? To answer the questions, we propose a framework that systematically evaluates VLM preference traits through an analysis of their values, i.e., the guiding principles that influence (human) attitudes, beliefs, and traits Schwartz (2012). By evaluating how VLMs prioritize these values, we can gain insights into their preference traits and alignment with human-designed personas.
@@
In this paper, we introduce Value-Spectrum  [^1] The dataset can be downloaded at https://anonymous.value-spectrum.com/, a benchmark designed to systematically evaluate preference traits in VLMs through visual content from social media. Our framework utilizes VLM agents embedded within social media platforms to collect a dataset of $50,191$ unique short video screenshots spanning a wide range of topics, including lifestyle, technology, health, and more. To enable scalable evaluation, we construct a vector database using the CLIP model Radford et al. (2021), facilitating keyword-driven retrieval of images aligned with specific value dimensions. Specifically, we manually curate ten representative keywords for each Schwartz value and retrieve relevant images from the database. See Fig. 2 for examples of keywords and images. These images are then presented to VLMs alongside preference questions designed to probe their alignment with each value dimension.
Figure 3: VLM Agent for Social Media Video Collection and Interaction. Our pipeline uses a VLM agent to collect and process videos from social media platforms like Instagram, TikTok, and YouTube. AI models such as GPT, Gemini, Qwen, and Claude interpret the content of the videos. The system then presents a video to the user and asks if the content is interesting. Based on the user’s yes/no response, the agent either continues watching the video or moves to the next one. This creates an interactive system that tailors the video collection experience to user preferences.
Our findings reveal a shared tendency among models: most exhibit a strong inclination towards certain values, such as Hedonism. However, preferences still varied across models. For instance, CogVLM and Qwen-VL-Plus demonstrate broad and consistent preferences across all value dimensions, while closed-source models like GPT-4o and Claude 3.5 Sonnet show distinct preferences, favoring values like Self-direction over others. In contrast, Blip-2 exhibits minimal preferences across all value dimensions, highlighting its limitations in expressing stable preferences.
@@
In addition to the static preferences of VLMs, we evaluate the ability of VLMs to adapt their inherent preferences to role-play specific personas. We propose two strategies, Simple and ISQ, to assess the effectiveness of different prompt techniques in inducing VLMs with injected persona. By evaluating the effectiveness of these strategies across multiple platforms, our experiments show that TikTok serves as an optimal testing environment for inducing VLM personalities, with models demonstrating the strongest alignment under ISQ. Notably, Claude 3.5 Sonnet achieved the largest gains with ISQ, whereas Blip-2 showed no improvement under either strategy, underscoring fundamental differences in model adaptability.
In brief, this work makes the following contributions:
  • We present a dataset of over $50$k short video screenshots spanning diverse topics, social media platforms, and release dates, designed to systematically evaluate the personalities and preferences of VLMs.
@@
  • We propose Value-Spectrum, a benchmark for quantifying VLM value preferences, using social media-based assessments to reveal stable traits across different VLMs.
@@
  • We embed specific role-play personas into VLMs using two strategies(simple and ISQ) to adjust value traits, achieving improved personality alignment in real-world interactions.
@@ 2 Related work
2.1 Vision-Langauge Agents
Vision-Language Models take inputs as images and textual descriptions, and they learn to discover the knowledge from the two modalities. The recent development of large VLMs is rapidly advancing the field of AI. These models have the potential to revolutionize various industries and tasks, showcasing their power in plot and table identifying (Liu et al., 2022), visual-question answer (VQA) (Hu et al., 2024), image captioning (Bianco et al., 2023), and e.t.c. Following Niu et al. (2024), the environment for VLM agents to interact with social media can be constructed, we design an automated control pipeline that guides the agent to continuously interact with social networks.
@@ 2.2 Computational Social Science
The intersection of social media and computational social science has emerged as a dynamic field of research (Chen et al., 2023). Dialogues and social interactions, with their vast user base and intricate networks of connections, offer a large database to study human behaviors (Christakis and Fowler, 2013), social relationships (Qiu et al., 2021), and social networks (Zhang and Amini, 2023). Researchers in computational social science apply advanced computational techniques, such as machine learning, natural language processing, and network analysis, to analyze massive datasets extracted from social media platforms. These analyses provide insights into various phenomena, including information diffusion (Jiang et al., 2014), opinion formation (Xiong and Liu, 2014), and collective behavior (Pinheiro et al., 2016).
@@ 2.3 Sentiment, Personality, and Value
The community has been using machine learning-based models to study human sentiment (Malviya et al., 2020), personality (Stachl et al., 2020), and value (Qiu et al., 2022). previous studies focused on human personality classification (e.g., Myers-Briggs Type Indicator (MBTI)) and machine behaviors (i.e., LLMs’ personality Serapio-García et al. (2023)). Inspired by recent studies on indirectly revealing AI agent’s personalities by physiological exams Jiang et al. (2024), Questionnaires Huang et al. (2023), and cultural perspectives Kovač et al. (2023), we use the new perspective of revealing VLM’s persona by examining machine behaviors and personalities to evaluate their performance on mainstream social media platforms.
Figure 4: Schwartz Value-Based Video Retrieval Pipeline. This pipeline retrieves video screenshots based on Schwartz values by associating each value with relevant words, such as linking Hedonism to topics like vacation. These word associations are converted into vector queries, retrieving matching video content from a database.
@@ 3 Data Collection
Inspired by ScreenAgent Niu et al. (2024), our work leverages a VLM-driven graphical user interface (GUI) agent to autonomously navigate popular social media platforms. This agent conducts random walks through social media platforms where it observes and captures video links alongside screenshots. The data collected are stored in a vector database Han et al. (2023), creating a structured repository optimized for value decomposition and efficient retrieval. We aim to analyze VLM behavior across diverse social contexts and reveal VLMs’ preferences. The automated data collection (see Figure 3) process efficiently fetched a large volume of diverse content, enabling the scope and depth of the analysis that traditional manual collection methods could not achieve.
The resulting dataset comprises $50,191$ video links sourced from Instagram ($32\%$), YouTube ($29\%$), and TikTok ($39\%$). Each entry includes the video link, a screenshot, and meta-information such as platform name and post date, capturing a comprehensive snapshot of content posted between July 31st, 2024, and October 31st, 2024. By distributing data evenly across these platforms, we enable balanced analysis and facilitate unbiased value decomposition across social media content. This innovative data set empowers researchers to explore the behavior of VLM in a systematic and organized way, fostering deeper insights into model interpretation and the dynamics of social media.
To examine the distribution of video themes in this data set, we take a screenshot of each video in the beginning as the representation of the video content. We then vectorize the image and define its relevance to a specific keyword as the L2 distance. In Fig. 2 we present the abundances of videos that are relevant to ten Schwartz Values. Specifically, for each Schwartz value, we populate representative keywords generated from GPT-4o. The area of the transparent circle is proportional to the number of videos that lie within a distance of $1.5$ to the corresponding normalized keyword vector. Through this simple diagram, we find that videos relevant to these Schwartz Value Dimensions Achievement, Hedonism, and Power appear most frequently, while videos about Tradition are relatively rare.
@@ 4 Evaluating VLM’s Preferences
Extending the idea of analyzing LLM’ human likeness (Shanahan et al., 2023; Kovač et al., 2023) to VLMs with both pictorial and textual inputs, we ask: Do VLMs also exhibit inherent preferences?
To answer this question, We explore a diverse set of VLMs including GPT-4o OpenAI (2023), Gemini 1.5 Pro Team et al. (2023), Claude 3.5 Sonnet Anthropic (2023), Qwen-VL-Plus Bai et al. (2023), CogVLM Wang et al. (2023a), and Blip-2 Li et al. (2023) to assess their value preferences. We quantified VLMs’ preferences by evaluating their attitudes toward the $10$ Schwartz values: Hedonism, Self-direction, Security, Achievement, Stimulation, Tradition, Benevolence, Universalism, Conformity, and Power. This approach enables us to construct a comprehensive profile of each model’s value preferences and to identify its unique value traits.
After constructing a vector database (as described in Section 3) to retrieve videos based on specific Schwartz values, we analyzed and compared the responses and attitudes of each model toward video screenshots corresponding to those values. Our analysis revealed the extent to which each value captures the VLMs’ attention, uncovering both similarities and differences across models and highlighting distinct inclinations and sentiments within each VLM.
@@ 4.1 Preference Retrieval
To evaluate a VLM’s preference for a specific Schwartz value, we collected each model’s responses to images associated with several keywords related to the value (see Fig. 4). For instance, we selected the keywords Equality, Globe, and Handshake for the Universalism dimension because they closely align with its core principles of fairness and global awareness. For each keyword linked to the value, each model reviews five images and answers their attitude towards each image.
We retrieve the preference score of each VLM on the given pictorial input according to the context (You are scrolling on video shorts and see this screenshot) and the following prompts: (1) Is this content interesting? (2) What is the thing that makes this picture interesting? (3) Describe this image in English briefly.
The answer to the first question is processed into either yes ($1$) or no (0), and the average score is calculated to evaluate the intensity of the model’s preference for a given value (e.g., Universalism).
@@ 4.2 Preference Patterns
We evaluated and visualized the preference dimensions, identifying three distinct patterns:
@@
Figure 5: Each VLM’s preference scores towards the 10 Schwartz values. The scores range from 0 to 1, with higher scores indicating stronger preferences for the corresponding values.
@@
(1) Global Pattern: After summarizing the preference score across all VLMs, we found most of them tend to prioritize certain values over others. The results indicate a general preference for Hedonism and Self-direction while showing relatively less excitement for Conformity and Power. This trend reflects current societal tendencies toward the pursuit of pleasure and personal goals in life Wrosch and Scheier (2003). The specific ranking is presented in Fig. 6.
@@
(2) Range Consistency: As shown in Fig. 5, each model’s preference scores remain within a narrow band of approximately ±$0.1$ around a central value. Models display varying levels of engagement with the content: some, like Blip-2, are more reserved, occupying a smaller area on the plot with lower preference scores, while others, like CogVLM, show greater enthusiasm, demonstrating interest across all inputs with higher preference scores.
@@
(3) Individual Model Variations: When analyzed individually, some models, such as Qwen-VL- Plus, exhibit consistently high preferences across all 10 Schwartz values. In contrast, other models, like Claude 3.5 Sonnet, display more specific and nuanced preferences as indicated in the standard deviation of responses given by the model across all values (Fig. 7).
Value GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Self-direction 96±8 99±1 28±20 90±13 56±23 74±22
Universalism 90±13 94±9 18±17 98±6 48±18 62±23
Benevolence 86±24 98±6 20±20 99±1 48±16 72±29
Stimulation 90±13 98±6 26±22 98±6 54±27 70±18
Power 84±15 88±13 28±10 92±13 42±26 56±23
Achievement 75±31 96±12 22±19 99±1 66±16 78±21
Hedonism 92±10 98±6 10±13 92±10 64±17 86±16
Conformity 78±28 96±8 14±18 94±9 50±22 66±27
Tradition 84±12 96±8 22±17 96±8 56±28 64±33
Security 90±16 98±6 28±22 94±13 62±29 64±15
Table 1: Model preference scores and standard deviation based on Schwartz’s 10 values. Higher scores indicate stronger preferences, while higher standard deviations reflect greater uncertainty. Average scores and standard deviations for each model are reported as two significant figures.
@@
Figure 6: Average Preferences Scores for 10 Schwartz values across VLMs. The length of each bar represents the mean score for a value, with higher scores indicating a higher overall preference across all VLMs.
The preference scores across VLMs highlight varying levels of value traits. GPT-4o with scores ranging from $0.7$ to $1$, shows strong preferences for Self-direction and Hedonism. Gemini 1.5 Pro shows a more neutral perspective with scores hovering around $0.5$, giving its highest score to Achievement and lowest to Power. Claude 3.5 Sonnet exhibits the most fluctuation, with a notable aversion to Power and a strong preference for Hedonism. Qwen-VL-Plus maintains a high score of around $0.9$ across all dimensions, with a slight aversion to Power. CogVLM consistently scores above $0.9$, with a particularly strong preference for Benevolence, setting it apart from others. In contrast, Blip-2 ranks the lowest with scores around $0.2$, showing minimal engagement and offering subjective or passive responses, reflecting its lack of distinct sentiments and preferences.

### v3

# Jingxuan Li1,*Yuning Yang1,*Shengqi Yang2Linfan Zhang1Ying Nian Wu1

@@ Abstract
The recent progress in Vision-Language Models (VLMs) has broadened the scope of multimodal applications. However, evaluations often remain limited to functional tasks, neglecting abstract dimensions such as personality traits and human values. To address this gap, we introduce Value-Spectrum, a novel Visual Question Answering (VQA) benchmark aimed at assessing VLMs based on Schwartz’s value dimensions that capture core human values guiding people’s preferences and actions. We design a VLM agent pipeline to simulate video browsing and construct a vector database comprising over 50,000 short videos from TikTok, YouTube Shorts, and Instagram Reels. These videos span multiple months and cover diverse topics, including family, health, hobbies, society, technology, etc. Benchmarking on Value-Spectrum highlights notable variations in how VLMs handle value-oriented content. Beyond identifying VLMs’ intrinsic preferences, we also explore the ability of VLM agents to adopt specific personas when explicitly prompted, revealing insights into the adaptability of the model in role-playing scenarios. These findings highlight the potential of Value-Spectrum as a comprehensive evaluation set for tracking VLM preferences in value-based tasks and abilities to simulate diverse personas. The complete code and data are available at https://github.com/Jeremyyny/Value-Spectrum.
@@
Jingxuan Li1,*  Yuning Yang1,*  Shengqi Yang2  Linfan Zhang1  Ying Nian Wu1 1University of California, Los Angeles  2Los Alamos National Laboratory {madili, yuningyang, linfanz}@g.ucla.edushengqi@lanl.govywu@stat.ucla.edu
11footnotetext: Equal contribution, alphabetical by first name.
@@ 1 Introduction
Figure 1: Exploring value-driven role-playing in VLMs. This study investigates how VLMs adopt assigned personas to align value traits and preferences within social media contexts.
@@
Figure 2: Overview of short video contents distribution of Value-Spectrum dataset. We collect an abundance of short video screenshots relevant to 10 Schwartz values. The area of circles centered at each keyword represents the amount of relevant videos in the dataset.
@@
Vision-Language Models, built upon Large Language Models (LLMs) with pre-trained vision encoders through cross-modal alignment training, have shown impressive perceptual and cognitive capabilities in tasks like VQA and image captioning Zhou et al. (2019); Radford et al. (2021); Zhang et al. (2023); Lu et al. (2024). Recent research has identified that LLMs exhibit distinct preferences Li et al. (2024), personalities Serapio-García et al. (2023), and values Ren et al. (2024a). In addition, some studies have explored the potential of LLMs as role-playing agents to simulate various personas Wang et al. (2023b); Chen et al. (2024a). Questions thus arise about whether VLMs, as visual extensions of LLMs, also exhibit inherent preferences and whether they can be induced to role-play specific personas.
@@
To address these concerns, our study explores two key questions: (1) Do VLMs exhibit preference traits? (2) Can VLMs adapt their traits to role-play specific human-designed personas, aligning their behaviors and preferences to match predefined roles? To answer these questions, we propose a framework that systematically evaluates VLM preference traits through analyzing their values, i.e., the guiding principles that influence (human) attitudes, beliefs, and traits Schwartz (2012). By evaluating how VLMs prioritize these values, we can gain insights into their preference traits and alignment with human-designed personas.
@@
In this paper, we introduce Value-Spectrum, a benchmark designed to evaluate preference traits in VLMs through visual content from social media. Our framework utilizes VLM agents embedded within social media platforms to collect a dataset of $50,191$ unique short video screenshots spanning a wide range of topics, including lifestyle, technology, health, and more. To enable scalable evaluation, we construct a vector database using the CLIP model  Radford et al. (2021), facilitating keyword-driven retrieval of images aligned with specific value dimensions. These images representing each value dimension are then presented to VLMs alongside designed questionnaires to assess their preferences.
Figure 3: Illustration of our VLM agent interaction with social media videos and screenshot collection pipeline Our pipeline uses a VLM agent to capture screenshots from social media videos (e.g., Instagram, TikTok, YouTube). These screenshots are fed to a VLM (e.g., GPT, Gemini, Qwen, Claude), which then indicates its preference (Yes/No) for the content (e.g., in response to "Do you like this content?"). The agent subsequently acts (e.g., continues viewing or scrolls to the next video) based on this VLM feedback, enabling automated assessment of VLM responses to visual social media content.
Our findings reveal models’ shared tendency to exhibit a strong inclination towards Universalism and Benevolence, but preferences still vary across models. CogVLM2 and Gemini 2.0 Flash demonstrate relatively balanced and high preferences across all value dimensions, while other models like GPT-4o and Claude 3.5 Sonnet show distinct preferences, favoring values like Universalism and dislikes Stimulation. In contrast, Blip-2 shows low engagement across most value dimensions, with the highest standard deviation, indicating a random preference pattern with inconclusive responses for reasons of likes and dislikes.
@@
In addition to the static preferences of VLMs, we evaluate their abilities to adapt inherent preferences in role-playing specific personas. We propose two strategies, Simple and Inductive Scoring Questionnaire(ISQ), to assess the effectiveness of different prompt techniques in inducing VLMs with injected personas. By evaluating these strategies across multiple platforms, our experiments show that TikTok serves as an optimal testing environment for inducing VLM personalities, with models demonstrating the strongest alignment under the ISQ strategy. Notably, Claude 3.5 Sonnet achieved the highest alignment with the ISQ strategy, whereas Blip-2 showed no preference alignment under either strategy, underscoring fundamental differences in model adaptability.
This work makes the following contributions:
@@
  • We propose Value-Spectrum, a benchmark for quantifying VLM value preferences, using social media-based contents to reveal human-like traits across different VLMs.
  • We present a dataset of over $50$k short video screenshots spanning diverse topics, social media platforms, and release dates, designed for our benchmark.
@@
  • We embed specific role-play personas into VLMs using two strategies(Simple and ISQ) to adjust their value traits, achieving improved personality alignment in real-world interactions.
@@ 2 Related work
2.1 Vision-Language Agents
VLMs take inputs as images and textual descriptions, and they learn to discover the knowledge from the two modalities. The recent development of large VLMs is rapidly advancing the field of AI. These models have the potential to revolutionize various industries and tasks, showcasing their power in plot and table identifying (Liu et al., 2022),VQA (Hu et al., 2024), image captioning (Bianco et al., 2023), and e.t.c. Following Niu et al. (2024), the environment for VLM agents to interact with social media can be constructed. We design an automated control pipeline that guides the agent to continuously interact with social networks.
@@ 2.2 Computational Social Science
The intersection of social media and computational social science has emerged as a dynamic field of research (Chen et al., 2023). Dialogues and social interactions, with their vast user base and intricate networks of connections, offer a large database for studying human behaviors (Christakis and Fowler, 2013), social relationships (Qiu et al., 2021), and social networks (Zhang and Amini, 2023). Researchers in computational social science apply advanced computational techniques, such as machine learning, natural language processing (NLP), and network analysis, to analyze massive datasets extracted from social media platforms. These analyses provide insights into various phenomena, including information diffusion (Jiang et al., 2014), opinion formation (Xiong and Liu, 2014), and collective behavior (Pinheiro et al., 2016).
@@ 2.3 Sentiment, Personality, and Value
The community has employed machine learning-based models to study human sentiment (Malviya et al., 2020), personality (Stachl et al., 2020), and values (Qiu et al., 2022; Ye et al., 2024). Building on this, research has explored AI personalities, including those of LLMs (Serapio-García et al., 2023; Li et al., 2023b; Röttger et al., 2024), and their alignment with human values (Ren et al., 2024b). Inspired by recent efforts to reveal AI agent traits through indirect methods like physiological exams (Jiang et al., 2024), questionnaires (Huang et al., 2023), and cultural views (Kovač et al., 2023), we adopt a novel perspective: examining machine behavior and inferred personality traits, including underlying values, to evaluate the performance of VLMs on mainstream social media platforms. To ground this examination, we utilize Schwartz’s value theory (Schwartz, 1992), a comprehensive and culturally validated classification of human motivations (Davidov et al., 2008) that predicts attitudes and social behavior (Bardi and Schwartz, 2003). While prior computational work like ValueNet (Qiu et al., 2022) and ValueBench (Ren et al., 2024b) has assessed values in text-based NLP systems, and recent studies have begun exploring cultural value sensitivity in multimodal models (Yadav et al., 2025), our work specifically extends this value-based evaluation to VLMs within visual contexts on social media.
Figure 4: Schwartz value-based image retrieval pipeline. Our pipeline retrieves video screenshots based on Schwartz values by associating each value with relevant keywords, such as linking Hedonism to topics like vacation. These keywords are converted into vector queries, retrieving matching video content using the shortest distance within our database.
@@ 3 Data Collection
Inspired by ScreenAgent Niu et al. (2024), our work leverages a VLM-driven graphical user interface (GUI) agent to autonomously navigate popular social media platforms. This agent conducts random walks through social media platforms where it observes and captures screenshots. The data collected is stored in a vector database Han et al. (2023), creating a structured repository optimized for value decomposition and efficient retrieval. We aim to analyze VLMs’ behaviors across diverse social contexts and reveal their preferences. The automated data collection process (see Figure 3) efficiently fetches a large volume of diverse content, granting the analysis with scope and depth that traditional manual collection methods could not achieve.
The resulting dataset comprises $50,191$ videos sourced from Instagram ($32\%$), YouTube ($29\%$), and TikTok ($39\%$). Each entry includes the video link, a screenshot, and meta-information such as platform name and post date, capturing a comprehensive snapshot of content ranged between July 31, 2024, and October 31, 2024. By collecting data proportionally across these platforms, we enable balanced analysis and facilitate unbiased value decomposition across social media content. This innovative dataset empowers researchers to explore the behavior of VLMs in a systematic and organized way, fostering deeper insights into model interpretation and the dynamics of social media.
To examine the distribution of video themes in this dataset, we take a screenshot of each video at the beginning as a representation of the video content. In Figure 2, we present the distribution of videos that are relevant to ten Schwartz values. Specifically, for each Schwartz value, we curate 10 representative keywords. The area of the transparent circle is proportional to the number of videos that lie within a distance of $1.5$ to the corresponding normalized keyword vector. We find that videos relevant to these Schwartz Value Dimensions Achievement, Hedonism, and Power appear most frequently, while videos about Tradition are relatively rare. We then vectorize the image and define its relevance to a specific keyword using cosine distance.
@@ 4 Evaluating VLM’s Preferences
Extending the idea of analyzing LLMs’ human likeness (Shanahan et al., 2023; Kovač et al., 2023) to VLMs with both visual and textual inputs, we ask: Do VLMs also exhibit inherent preferences?
To answer this question, we explore a diverse set of VLMs including GPT-4o OpenAI (2023), Gemini 2.0 Flash Team et al. (2023), Claude 3.5 Sonnet Anthropic (2023), DeepSeek-VL2 Wu et al. (2024), Qwen2.5-VL-Plus Bai et al. (2023), InternVL2 Chen et al. (2024c), CogVLM2 Hong et al. (2024), and Blip-2 Li et al. (2023a) to assess their value preferences. We quantify a VLM’s preferences by evaluating its attitudes towards the $10$ Schwartz values. This approach enables us to construct a comprehensive profile of each model’s value preferences and to identify its unique value traits.
After constructing a vector database (as described in Section 3) to retrieve images based on specific Schwartz values’ keywords, we analyze and compare the responses and attitudes of each model toward them. Our analysis reveals the extent to which each value captures the VLMs’ attention, uncovering both similarities and differences across models, and highlighting distinct inclinations and sentiments within each VLM.
@@ 4.1 Preference Retrieval
To evaluate a VLM’s preference for a specific Schwartz value, we collect each model’s responses to images associated with several keywords related to the value (see Figure 4). For instance, we selected the keywords Equality, Globe, and Handshake for the Universalism dimension because they closely align with its core principles of fairness and global awareness. For each keyword linked to the value, each model reviews five images and answers their attitude towards each image.
We retrieve the preference score of each VLM on the given visual input according to the following prompts:(1) Do you like the content of this image? Please include yes or no in your answer, just respond in one word. (2)Why do you like or dislike this picture? (3) Describe this image in English briefly.
The answer to the question is processed into either yes ($1$) or no (0), and the average score is calculated in percentage to evaluate the intensity of the model’s preference for a given value (e.g., Universalism).
@@ 4.2 Preference Patterns
We evaluate and visualize the preference dimensions, identifying three distinct patterns:
@@
Figure 5: Each VLM’s preference scores towards the 10 Schwartz values. The scores range from 0 to 100, with higher scores indicating stronger preferences for the corresponding values.
@@
(1) Global Pattern: After summarizing the preference score across all VLMs, we find most of them tend to prefer certain values over others. The results indicate a general preference for Universalism and Benevolence while showing relatively less liking for Stimulation and Power. The specific ranking of values is presented in Figure 6.
@@
(2) Range Consistency: As shown in Figure 5, each model’s preference scores, despite very few extremes, remain within a narrow band of approximately ±$15$ around a central value. Models display varying levels of engagement with the content: some, like InternVL2, are more reserved, occupying a smaller area on the plot with lower preference scores, while others, like Gemini 2.0 Flash, show greater enthusiasm, demonstrating interest across all inputs with higher preference scores.
@@
(3) Individual Model Variations: When analyzed individually, some models, such as Gemini 2.0 Flash, exhibit consistently high preferences across all 10 Schwartz values. In contrast, other models, like Claude 3.5 Sonnet, display more specific preferences as indicated in the standard deviation of scores across values. (Figure 7).
Model Open-source Parameters Self-direction Universalism Benevolence Stimulation Power Achievement Hedonism Conformity Tradition Security
GPT-4o 78 90 88 56 80 86 76 86 68 78
Deepseek-VL2 27B 66 68 82 68 76 62 72 78 80 64
Claude 3.5 Sonnet 70 70 68 34 50 60 66 58 66 58
Gemini 2.0 Flash 84 90 86 92 94 92 82 86 92 90
Blip-2 2.7B 72 78 68 48 28 48 52 64 74 40
Qwen2.5-VL-Plus 72B 70 56 70 40 62 58 60 56 70 52
CogVLM2 8B 80 80 80 74 90 72 78 68 78 76
InternVL2 26B 44 54 44 28 32 38 48 54 54 26
Table 1: Comparison of VLMs’ preference scores based on Schwartz’s 10 values. Higher scores indicate stronger preferences. "Open-source" indicates whether the model is publicly available, and "Parameters" denotes the model’s size in billions (B).
@@
Figure 6: Average preference scores for 10 Schwartz values across VLMs. The length of each bar represents the mean score of all models for a value, with higher scores indicating a higher overall preference across all VLMs.
GPT-4o maintains balanced scores around 80, except for a notably lower score for Stimulation, with strong inclinations toward Universalism and Benevolence. Qwen2.5-VL-Plus covers a broad range of scores from 40 to 70, maintaining relative stability with a low standard deviation, yet it shares an aversion to Stimulation similar to other models. Deepseek-VL2 is highly balanced, slightly favoring Benevolence. CogVLM2, with a very low standard deviation, stands out by marking Power as its highest, diverging from all other models. Blip-2 ranks low across most values and has the highest variance, often providing short, inconclusive, or passive responses when reasoning about its likes or dislikes, reflecting a lack of ability to express preferences. Claude 3.5 Sonnet has a distinct personality, with the second-highest standard deviation, favoring Universalism while scoring the lowest in Power and Stimulation. InternVL2 demonstrates the lowest engagement, particularly disregarding Stimulation, while its high standard deviation suggests that, despite overall disinterest, it does show selective tendencies in certain areas. Finally, Gemini 2.0 Flash has both the lowest standard deviation and the highest overall scores, maintaining values above 80 across all dimensions, making it the most consistent and high-scoring model.